OpenStudy (anonymous):

Solve 7x + 18 = 4 for x.

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

x= -2 7 x -2 = -14 + 18 = 4

whitemonsterbunny17 (whitemonsterbunny17):

@-the-humdrum- don't forget to explain your answer.. And next time, guide the user step by step on how to solve the problem.. Giving out direct answers is against this site's CoC (Code of Conduct).